OpenStudy (henrietepurina):

PLEASE HELP ON ALGEBRA 1 QUESTION Don't just give answer.. explain plz :) Medal + Fan

OpenStudy (anonymous):

f(x)=(x+3)(x+2)(x-2) \[=(x+3)(x^2-4)=?\] simplify it.

OpenStudy (henrietepurina):

will do @surjithayer x^3+3x^2−4x−12

OpenStudy (henrietepurina):

click refresh to remove those question marks :)

OpenStudy (anonymous):

correct

OpenStudy (henrietepurina):

what about that equation?

OpenStudy (henrietepurina):

Oh wait... that's A right?

OpenStudy (henrietepurina):

@surjithayer

OpenStudy (anonymous):

correct.

OpenStudy (henrietepurina):

Thanks so @surjithayer :) I got a 50/50 thanks to you :)

OpenStudy (anonymous):

you are welcome.